OpenStudy (anonymous):

What is the sum of the finite arithmetic series? 26 + 29 + 32 + 35 + 38 + 41 + 44

OpenStudy (anonymous):

\[3 * 70 + 35 = Your Answer\]

OpenStudy (anonymous):

Where did those numbers come from @Light&Happiness ?

OpenStudy (anonymous):

Well, you must pay attention at first and last terms, sum of them is 70. second term and previous term from the last , sum of them is 70, too,. next is the same, and you have the middle term . 3*70 + middle term to get the answer! (Which is 35.)

OpenStudy (anonymous):

3∗70+35 = 245 right?

OpenStudy (anonymous):

3∗70+35 = 245 right?

OpenStudy (anonymous):

Yep! That's right. That is your answer.